# GCL-OT: Graph Contrastive Learning with Optimal Transport for Heterophilic Text-Attributed Graphs

### v1

@@ Abstract
Recently, structure–text contrastive learning has shown promising performance on text-attributed graphs by leveraging the complementary strengths of graph neural networks and language models. However, existing methods typically rely on homophily assumptions in similarity estimation and hard optimization objectives, which limit their applicability to heterophilic graphs. Although existing methods can mitigate heterophily through structural adjustments or neighbor aggregation, they usually treat textual embeddings as static targets, leading to suboptimal alignment. In this work, we identify the multi-granular heterophily in text-attributed graphs, including complete heterophily, partial heterophily, and latent homophily, which makes structure–text alignment particularly challenging due to mixed, noisy, and missing semantic correlations. To achieve flexible and bidirectional alignment, we propose GCL-OT, a novel graph contrastive learning framework with optimal transport, equipped with tailored mechanisms for each type of heterophily. Specifically, for partial heterophily, we design a RealSoftMax-based similarity estimator to emphasize key neighbor-word interactions while easing background noise. For complete heterophily, we introduce a prompt-based filter that adaptively excludes irrelevant noise during optimal transport alignment. Furthermore, we incorporate OT-guided soft supervision to uncover potential neighbors with similar semantics, enhancing the learning of latent homophily. Theoretical analysis shows that GCL-OT can improve the mutual information bound and Bayes error guarantees. Extensive experiments on nine benchmarks show that GCL-OT consistently outperforms state-of-the-art methods, verifying its effectiveness and robustness.
@@ Introduction
Figure 1: Examples and empirical analysis of multi-granular heterophily in TAGs. Node colors denote categories. $H_{N}/H_{E}$: node/edge heterophily, $R_{\text{NTD}}/R_{\text{NWD}}$: neighbor token/sentence dissimilarity, $R_{\text{UTS}}$: similarity of unconnected nodes.
@@
Text-attributed graphs (TAGs) represent text entities as nodes and their relationships as edges, widely used across diverse real-world domains, such as academic citations, web hyperlinks, and e-commerce recommendations (yan2023tagbenchmark). Recent advances in Large Language Models (LLMs) have shown substantial capability to capture the semantic richness, prompting researchers to combine Language Models (LMs) with Graph Neural Networks (GNNs) on TAGs for performance enhancement (he2024TAPE; pan2024Integrating; Liu2025GFM). Among them, joint training strategies using Graph Contrastive Learning (GCL) (brannon2024congrat; Fang2024GAugLLM) enhance alignment between textual and structural representations by optimizing mutual information to distinguish positive and negative examples.
@@
However, heterophilic TAGs are prevalent in reality due to the principle of opposites attracting, e.g., dating networks. Early GNN and GCL methods typically adopt the non-local neighbor extension and architectural refinement strategies (chen2024polygcl; wang2024HeterGCL). Despite effectiveness, these approaches may be limited by the shallow embedding (Mikolov2013Distributed; zhang2010understanding). Few methods study heterophily by combining deep semantic representations from LMs. Only LLM4HeG (wu2024LLM4HeG) and LEMP4HG (wang2025LEMP4HG) explore utilizing LLMs/LMs for heterophilic edge discrimination and reweighting, but face sub-optimal challenges with sparse or noisy data in the cascading pipelines (pan2024Integrating; Liu2025GFM).
@@
In this paper, we identify that TAGs exhibit multi-granular heterophily from three node-centric perspectives grounded in the text-feature view, as illustrated in Figure 1(a): (i) Partial heterophily occurs when only part of a node’s text semantically aligns with its neighbors, and vice versa. Existing methods still struggle to capture subtle semantic mismatches (song2023orderedGNN). (ii) Complete heterophily appears when a node’s text has no relevance with its neighbors at all, typically caused by meaningless texts or connectivity (such as random co-purchases). Even state-of-the-art models can be misled by such signals (li2025HeTGB). (iii) Latent homophily denotes potential semantic neighbors that are disconnected, often due to missing or implicit links. While multi-hop propagation (zhu2020H2GCN) and potential neighbor discovery (Pei2020GeomGCN) exploit such signals, fully leveraging them remains open. As shown in Figure 1(b), these patterns are prevalent in real-world TAGs and naturally induce many-to-many (N:N) alignment between structural and textual neighborhoods, whereas DGIs nearly N:1 summary pairing and InfoNCEs approximately 1:1 correspondence fail to capture this. Fortunately, optimal transport (OT) can fractionally allocate mass for soft N:N alignment when 1:1 is ambiguous, enabling mutual coverage via OT constraints to avoid one-sided greedy matches(Xu2023GWF).
@@
To address the multi-granular heterophily challenges in TAGs, we further propose GCL-OT, a novel GCL framework with tailored mechanisms grounded in OT theory, aiming to hierarchically and softly align structural and textual representations. Specifically, for partial heterophily, the RealSoftMax-based similarity estimation can identify and emphasize the most relevant word–neighbor pairs bidirectionally. For complete heterophily, the global filter-prompt strategy is used to eliminate the negative influence of irrelevant embeddings during alignment. For latent homophily, the OT assignment is used as the auxiliary supervision within the GCL objective to find hidden neighbors. We analyze GCL-OT using mutual information (MI) theory, showing it can tighten the InfoNCE MI lower bound and reduce Bayes error in downstream node classification. Extensive experiments on nine TAG benchmarks demonstrate that GCL-OT consistently outperforms strong baselines, highlighting its robustness and potential in both homophilic and heterophilic settings. Our main contributions are summarized as follows:
@@
Representation learning on Text‑Attributed Graphs (TAGs) has attracted significant attention in graph machine learning. Traditional approaches typically couple shallow textual embeddings, such as bag-of-words or FastText, with GNNs (kipf2017GCN; hamilton2017SAGE; velivckovic2018GAT). Recent efforts combine GNNs with LMs for richer textual encoding in cascading, iterative, and parallel manners. Cascading methods freeze GNN or LM as an auxiliary to enrich the other, while the two-stage, separately encoding often produce sub-optimal integration (he2024TAPE; wu2024LLM4HeG; wang2025LEMP4HG). Iterative methods alternately optimize LMs and GNNs under the shared objective, improving label efficiency but potentially increasing computational costs (yang2021graphformers; zhao2023GLEM). Parallel methods, mostly based on GCL, align LMs and GNNs in a shared space (zhu2024ENGINE; Fang2024GAugLLM). Congrat (brannon2024congrat) and GraphGPT (tang2024graphgpt) align them by node-level InfoNCE objectives, which are intuitive but vulnerable to complex relations and noise. HASE-code (zhang2024hasecode) and G2P2 (wen2023G2P2) extend the alignment granularity to subgraphs, but may risk over-smoothing due to high-order structures. Few approaches consider heterophily in TAGs. LLM4HeG (wu2024LLM4HeG) uses LLMs to label heterophilic edges from node-pair texts for reweighting. LEMP4HG (wang2025LEMP4HG) refines it via key-pair selection and attention integration. However, their two-stage pipelines make them sensitive to noisy text (pan2024Integrating; Liu2025GFM).
@@
Heterophily is generally considered a key challenge for GNN performance. Current researches typically fall into non-local neighbor extension and architectural refinement strategies (zheng2022graph; gong2024towards). The first extends non-local neighbors with similar attributes through mixing high-order neighbors (abu2019mixhop; song2023orderedGNN) or discovering potential neighbors based on various distance (Pei2020GeomGCN). The second refines GNN architecture by enhancing message aggregation from similar neighbors while minimizing the influence of dissimilar ones (bo2021beyond; zhu2021graph; liang2023predicting), or mixing layers to capture information from different neighbor ranges (xu2018representation; zhu2020H2GCN). Among them, recent GCL methods offer noteworthy advancements (chen2025m3pgcl). PolyGCL (chen2024polygcl) designs and contrasts learnable spectral polynomial filters for varying homophily levels, while HeterGCL (wang2024HeterGCL) incorporates structural and semantic modules to utilize label-inconsistent signals effectively. However, these methods rely on static shallow embeddings and struggle to capture context-aware information and complex semantic relationships, limiting their effectiveness in leveraging text attributes (he2024TAPE).
Graph Contrastive Learning with OT.
Optimal transport (OT) (Monge1781; Villani2009OT) is a mathematical framework for measuring distances between distributions, finding the most cost-efficient way to transform one distribution into another (Vincent2022OTGNN; lin2023multi; Xu2023GWF). In Graph Contrastive Learning (GCL), traditional InfoNCE loss often relies on hard alignment with pairing positive and negative samples, which may introduce representation bias. To this end, researchers solve the problems caused by non-aligned contrast views or structural differences in the graph by introducing OT distance for soft alignment, subgraph generation, or cross-view comparison (wang2023galopa; deng2025THESAURUS), thereby enhance node representation learning, which subsequently benefits downstream tasks such as node clustering  (Zhang2024CSOT; Wang2024PVCCL; deng2025THESAURUS), node classification (xie2024SGEC; Amadou2025FOSSIL; Zhu2022RoSA; Liu2025STAR), and node anomaly detection (Wang2023ACT). Prior OT-based GCL methods (e.g., THESAURUS, FOSSIL) already treat OT as a soft-alignment remedy to InfoNCE’s hard pairing, but they operate on homogeneous graphs or prototype graphs.
@@
Raw textual attributes are typically noisy and semantically ambiguous. To enrich their expressiveness, we augment each node’s text using a task-specific prompt issued to a frozen large language model (LLM) following (he2024TAPE). The original and prompted texts are concatenated to form an enhanced textual description $t_{i}^{\text{aug}}$ for each node $v_{i}$.
Then, we encode $t_{i}^{\text{aug}}$ with a lightweight PLM (e.g., DistilBERT (sanh2019distilbert)) to obtain token-level embeddings $\mathbf{H}^{\varpi}_{i}\in\mathbb{R}^{W\times d_{t}}$ and the sentence-level embedding $\mathbf{h}^{t}_{i}\in\mathbb{R}^{d_{t}}$. In parallel, we encode structures using a graph neural network (GNN). Node features are initialized and propagated through the GNN to produce structural embeddings $\mathbf{H}^{\zeta}\in\mathbb{R}^{N\times d_{s}}$, where $d_{s}$ is the structural dimension. For each node $v_{i}$, we derive a neighborhood-level embedding matrix $\mathbf{H}^{\mathcal{N}}_{i}\in\mathbb{R}^{K\times d_{s}}$ by aggregating and linear projecting the embeddings of its neighborhood $\mathcal{N}(v_{i})=\{v_{k}\mid e_{ik}\in\mathcal{E}\}\cup\{v_{i}\}$, where $K=|\mathcal{N}(v_{i})|$. These four views provide the foundation for contrastive objectives.
@@
where $\mathbf{z}$ is a learnable vector. Embeddings whose maximum similarity falls below corresponding $\mathbf{z}$ are instead align with the prompt vector during the subsequent alignment.
@@
As illustrated in Equation (3), OT can realign each node’s structure or text to multiple related texts or structures based on similarity, effectively resolving the potential problem between structures and texts. Let the kernel matrix $\mathbf{K}=\exp(\bar{\mathbf{S}}/\varepsilon)$, and we approximate $\mathbf{K}$ with a non-negative low-rank factorization
@@
where $\mathbf{U}\in\mathbb{R}_{+}^{(N+1)\times r}$ and $\mathbf{V}\in\mathbb{R}_{+}^{(N+1)\times r}$ are non-negative basis matrices, $\operatorname{diag}(\Upsilon)\in\mathbb{R}^{r\times r}$ turns the weight vector into a diagonal matrix, and target rank $r\leq N$. The optimal $\mathbf{Q}^{*}$ of Equation (3) has a simple normalized exponential matrix solution by LRSinkhorn (Scetbon2021lrsinkhorn):
@@
Let $\mathcal{L}_{\text{InfoNCE}}$ denote the standard InfoNCE loss between structural embeddings $\mathbf{H}^{\zeta}$ and textual embeddings $\mathbf{H}^{t}$. According to the standard InfoNCE MI lower bound (oord2019infonce), $\mathcal{L}_{\text{LHM}}$ provides a tighter variational lower bound than InfoNCE
@@
According to the standard InfoNCE MI lower bound (oord2019infonce), the latent-homophily objective $\mathcal{L}_{\text{LHM}}$ also satisfies
@@
Optimizing objective in Equation (11), GCL-OT provides a tighter upper bound on the downstream Bayes error (tsai2021sslmultiview; xiao2022decoupled) compared to $\mathcal{L}_{\textbf{InfoNCE}}$ and $\mathcal{L}_{\textbf{NC}}$. This suggests that downstream tasks can benefit from the learned representations obtained through our loss.
@@
To evaluate the performance of GCL-OT, the node classification tasks are conducted on 9 real-world datasets across different homophily degrees. Cora, PubMed (sen2008collective_pubmed), ArXiv, and Products (hu2020ogb) generally regarded as homophilic, whereas Amazon (platonov2023critical), ArXiv23 (he2024TAPE), Wisconsin, Cornell, and Texas (Pei2020GeomGCN) fall into heterophilic category. Original textual information for heterophilic graphs was collected from publicly available sources (leskovec2007dynamics; craven1998learning; wu2024LLM4HeG).
Experiments are conducted on one NVIDIA 4090 GPU and three NVIDIA 3080 GPUs. Text augmentation uses GPT-3.5 following (he2024TAPE). Text encoder adopts partially frozen DistilBERT (sanh2019distilbert) with six layers. Graph encoder adopts GCN (kipf2017GCN), GAT (velivckovic2017graph_gat), and GraphSAGE (hamilton2017SAGE), each configured with two layers. GAT utilizes eight attention heads, and GraphSAGE samples all neighbors. The classifier adopts a three-layer MLP with ReLU. Dimension of Text and Graph encoder is set to 768, with learning rates of 0.01 and 2e-5, weight decay at 0.0005, dropout from 0 to 0.8, and a batch size of 512. ArXiv and Products follow the standard splits from (hu2020ogb), the other datasets use random splits (60/20/20) following (he2024TAPE). Baseline hyperparameters follow prior work. All experiments report the average accuracy and standard deviation over 10 random seeds.
@@
We first study the improvements over InfoNCE across various heterophily metrics by perturbing edges and texts. As show in Figure 3, GCL-OT achieves substantial gains over the InfoNCE variant, obtained by replacing $\mathcal{L}_{\text{GCL-OT}}$ with $\mathcal{L}_{\text{InfoNCE}}$, with significant improvements under mixed-label neighborhoods and strong semantic heterophily. These trends indicate that GCL-OT can handle label mixing and semantic disagreement within neighborhood, while the abundance of purely heterophilic neighborhoods and unlinked similar pairs may limit its benefits.
@@
(a) DistilBERT on text perturbation
(b) GCL-OT on text perturbation

### v2

@@ Abstract
Recently, structure–text contrastive learning has shown promising performance on text-attributed graphs by leveraging the complementary strengths of graph neural networks and language models. However, existing methods typically rely on homophily assumptions in similarity estimation and hard optimization objectives, which limit their applicability to heterophilic graphs. Although existing methods can mitigate heterophily through structural adjustments or neighbor aggregation, they usually treat textual embeddings as static targets, leading to suboptimal alignment. In this work, we identify the multi-granular heterophily in text-attributed graphs, including complete heterophily, partial heterophily, and latent homophily, which makes structure–text alignment particularly challenging due to mixed, noisy, and missing semantic correlations. To achieve flexible and bidirectional alignment, we propose GCL-OT, a novel graph contrastive learning framework with optimal transport, equipped with tailored mechanisms for each type of heterophily. Specifically, for partial heterophily, we design a RealSoftMax-based similarity estimator to emphasize key neighbor-word interactions while easing background noise. For complete heterophily, we introduce a prompt-based filter that adaptively excludes irrelevant noise during optimal transport alignment. Furthermore, we incorporate OT-guided soft supervision to uncover potential neighbors with similar semantics, enhancing the learning of latent homophily. Theoretical analysis shows that GCL-OT can improve the mutual information bound and Bayes error guarantees. Extensive experiments on nine benchmarks show that GCL-OT outperforms state-of-the-art methods, demonstrating its effectiveness and robustness.
@@ Introduction
Figure 1: Analysis of multi-granular heterophily in TAGs. Node colors denote categories. $H_{N}/H_{E}$: node/edge heterophily, $R_{\text{NTD}}/R_{\text{NWD}}$: neighbor token/sentence dissimilarity, $R_{\text{UTS}}$: similarity of unconnected nodes.
@@
Text-attributed graphs (TAGs) represent text entities as nodes and their relationships as edges, widely used across diverse real-world domains, such as academic citations, web hyperlinks, and e-commerce recommendations (Yan et al. 2023). Recent advances in Large Language Models (LLMs) have shown substantial capability to capture the semantic richness, prompting researchers to combine Language Models (LMs) with Graph Neural Networks (GNNs) on TAGs for performance enhancement (He et al. 2024; Pan et al. 2024; Liu et al. 2025a). Among them, joint training strategies using Graph Contrastive Learning (GCL) (Brannon et al. 2024; Fang et al. 2024) enhance alignment between textual and structural representations by optimizing mutual information to distinguish positive and negative examples.
@@
However, heterophilic TAGs are prevalent in reality due to the principle of opposites attracting, e.g., dating networks. Early GNN and GCL methods typically adopt the non-local neighbor extension and architectural refinement strategies (Chen et al. 2024; Wang et al. 2024a). Despite effectiveness, these approaches may be limited by the shallow embedding (Mikolov et al. 2013; Zhang et al. 2010). Few methods study heterophily by combining deep semantic representations from LMs. Only LLM4HeG (Wu et al. 2025) and LEMP4HG (Wang and Cheng 2025) explore utilizing LLMs/LMs for heterophilic edge discrimination and reweighting, but face sub-optimal challenges with sparse or noisy data in the cascading pipelines (Pan et al. 2024; Liu et al. 2025a).
@@
In this paper, we identify that TAGs exhibit multi-granular heterophily from three node-centric perspectives grounded in the text-feature view, as illustrated in Figure 1(a): (i) Partial heterophily occurs when only part of a node’s text semantically aligns with its neighbors, and vice versa. Existing methods still struggle to capture subtle semantic mismatches (Song et al. 2023). (ii) Complete heterophily appears when a node’s text has no relevance with its neighbors at all, typically caused by meaningless texts or connectivity (such as random co-purchases). Even state-of-the-art models can be misled by such signals (Li et al. 2025). (iii) Latent homophily denotes potential semantic neighbors that are disconnected, often due to missing or implicit links. While multi-hop propagation (Zhu et al. 2020a) and potential neighbor discovery (Pei et al. 2020) exploit such signals, fully leveraging them remains open. As shown in Figure 1(b), these patterns are prevalent in real-world TAGs and naturally induce many-to-many (N:N) alignment between structural and textual neighborhoods, whereas DGIs nearly N:1 summary pairing and InfoNCEs approximately 1:1 correspondence fail to capture this. Fortunately, optimal transport (OT) can fractionally allocate mass for soft N:N alignment when 1:1 is ambiguous, enabling mutual coverage via OT constraints to avoid one-sided greedy matches(Xu et al. 2023).
@@
To address the multi-granular heterophily challenges in TAGs, we further propose GCL-OT, a novel GCL framework with tailored mechanisms grounded in OT theory, aiming to hierarchically and softly align structural and textual representations. Specifically, for partial heterophily, the RealSoftMax-based similarity estimation can identify and emphasize the most relevant word–neighbor pairs bidirectionally. For complete heterophily, the global filter-prompt strategy is employed to mitigate the negative impact of irrelevant embeddings during alignment. For latent homophily, OT assignment serves as the auxiliary supervision within the GCL objective to discover hidden neighbors. We analyze GCL-OT using mutual information (MI) theory, showing it can tighten the InfoNCE MI lower bound and reduce Bayes error in downstream node classification. Extensive experiments on nine TAG benchmarks demonstrate that GCL-OT outperforms strong baselines, highlighting its robustness and potential in both homophilic and heterophilic settings. Our main contributions are summarized as follows:
@@
Representation learning on Text‑Attributed Graphs (TAGs) has attracted significant attention in graph machine learning. Traditional approaches typically couple shallow textual embeddings, such as bag-of-words or FastText, with GNNs (Kipf and Welling 2017; Hamilton et al. 2017; Veličković et al. 2018). Recent efforts combine GNNs with LMs for richer textual encoding in cascading, iterative, and parallel manners. Cascading methods freeze GNN or LM as an auxiliary to enrich the other, while the two-stage, separately encoding often produce sub-optimal integration (He et al. 2024; Wu et al. 2025; Wang and Cheng 2025). Iterative methods alternately optimize LMs and GNNs under the shared objective, improving label efficiency but potentially increasing computational costs (Yang et al. 2021; Zhao et al. 2023). Parallel methods, mostly based on GCL, align LMs and GNNs in a shared space (Zhu et al. 2024; Fang et al. 2024). Congrat (Brannon et al. 2024) and GraphGPT (Tang et al. 2024) align them by node-level InfoNCE objectives, which are intuitive but vulnerable to complex relations and noise. HASE-code (Zhang et al. 2024a) and G2P2 (Wen and Fang 2023) extend the alignment granularity to subgraphs, but may risk over-smoothing due to high-order structures. Few approaches consider heterophily in TAGs. LLM4HeG (Wu et al. 2025) uses LLMs to label heterophilic edges from node-pair texts for reweighting. LEMP4HG (Wang and Cheng 2025) refines it via key-pair selection and attention integration. However, their two-stage pipelines make them sensitive to noisy text (Pan et al. 2024; Liu et al. 2025a).
@@
Heterophily is generally considered a key challenge for GNN performance. Current researches typically fall into non-local neighbor extension and architectural refinement strategies (Zheng et al. 2022; Gong et al. 2024). The first extends non-local neighbors with similar attributes through mixing high-order neighbors (Abu-El-Haija et al. 2019; Song et al. 2023) or discovering potential neighbors based on various distances (Pei et al. 2020). The second refines GNN architecture by enhancing message aggregation from similar neighbors while minimizing the influence of dissimilar ones (Bo et al. 2021; Zhu et al. 2021; Liang et al. 2023), or mixing layers to capture information from different neighbor ranges (Xu et al. 2018; Zhu et al. 2020a). Among them, recent GCL methods offer noteworthy advancements (Chen et al. 2025). PolyGCL (Chen et al. 2024) designs and contrasts learnable spectral polynomial filters for varying homophily levels, while HeterGCL (Wang et al. 2024a) incorporates structural and semantic modules to utilize label-inconsistent signals effectively. However, these methods rely on static shallow embeddings and struggle to capture context-aware information and complex semantic relationships, limiting their effectiveness in leveraging text attributes (He et al. 2024).
@@
Graph Contrastive Learning with OT
Optimal transport (OT) (Monge 1781; Villani 2009) is a mathematical framework for measuring distances between distributions, finding the most cost-efficient way to transform one distribution into another (Vincent-Cuaz et al. 2022; Lin et al. 2023; Xu et al. 2023). In Graph Contrastive Learning (GCL), traditional InfoNCE loss often relies on hard alignment with pairing positive and negative samples, which may introduce representation bias. To this end, researchers solve the problems caused by non-aligned contrast views or structural differences in the graph by introducing OT distance for soft alignment, subgraph generation, or cross-view comparison (Wang et al. 2023b; Deng et al. 2025), thereby enhance node representation learning, which subsequently benefits downstream tasks such as node clustering  (Zhang et al. 2024b; Wang et al. 2024c; Deng et al. 2025), node classification (Xie and Giraldo 2024; SANGARE et al. 2025; Zhu et al. 2022; Liu et al. 2025b), and node anomaly detection (Wang et al. 2023a). Prior OT-based GCL methods (e.g., THESAURUS, FOSSIL) already treat OT as a soft-alignment remedy to InfoNCE’s hard pairing, but they operate on homogeneous graphs or prototype graphs.
@@
Raw textual attributes are typically noisy and semantically ambiguous. To enrich their expressiveness, we augment each node’s text using a task-specific prompt issued to a frozen large language model (LLM) following (He et al. 2024). The original and prompted texts are concatenated to form an enhanced textual description $t_{i}^{\text{aug}}$ for each node $v_{i}$.
Then, we encode $t_{i}^{\text{aug}}$ with a lightweight PLM (e.g., DistilBERT (Sanh et al. 2020)) to obtain token-level embeddings $\mathbf{H}^{\varpi}_{i}\in\mathbb{R}^{W\times d_{t}}$ and the sentence-level embedding $\mathbf{h}^{t}_{i}\in\mathbb{R}^{d_{t}}$. In parallel, we encode structures using a graph neural network (GNN). Node features are initialized and propagated through the GNN to produce structural embeddings $\mathbf{H}^{\zeta}\in\mathbb{R}^{N\times d_{s}}$, where $d_{s}$ is the structural dimension. For each node $v_{i}$, we derive a neighborhood-level embedding matrix $\mathbf{H}^{\mathcal{N}}_{i}\in\mathbb{R}^{K\times d_{s}}$ by aggregating and linear projecting the embeddings of its neighborhood $\mathcal{N}(v_{i})=\{v_{k}\mid e_{ik}\in\mathcal{E}\}\cup\{v_{i}\}$, where $K=|\mathcal{N}(v_{i})|$. These four views provide the foundation for contrastive objectives.
@@
where $\mathbf{z}$ is a learnable vector. Embeddings whose maximum similarity falls below the corresponding $\mathbf{z}$ are instead aligned with the prompt vector during the subsequent alignment.
@@
As shown in Equation (3), OT can realign each node’s structure or text with multiple related texts or structures based on similarity, effectively resolving potential issues between structures and texts. Let the kernel matrix $\mathbf{K}=\exp(\bar{\mathbf{S}}/\varepsilon)$, and we approximate $\mathbf{K}$ with a non-negative low-rank factorization
@@
where $\mathbf{U}\in\mathbb{R}_{+}^{(N+1)\times r}$ and $\mathbf{V}\in\mathbb{R}_{+}^{(N+1)\times r}$ are non-negative basis matrices, $\operatorname{diag}(\Upsilon)\in\mathbb{R}^{r\times r}$ turns the weight vector into a diagonal matrix, and target rank $r\leq N$. The optimal $\mathbf{Q}^{*}$ of Equation (3) has a simple normalized exponential matrix solution by LRSinkhorn (Scetbon et al. 2021):
@@
Let $\mathcal{L}_{\text{InfoNCE}}$ denote the standard InfoNCE loss between structural embeddings $\mathbf{H}^{\zeta}$ and textual embeddings $\mathbf{H}^{t}$. According to the standard InfoNCE MI lower bound (van den Oord et al. 2019), $\mathcal{L}_{\text{LHM}}$ provides a tighter variational lower bound than InfoNCE
@@
According to the standard InfoNCE MI lower bound (van den Oord et al. 2019), the latent-homophily objective $\mathcal{L}_{\text{LHM}}$ also satisfies
@@
Optimizing objective in Equation (11), GCL-OT provides a tighter upper bound on the downstream Bayes error (Tsai et al. 2021; Xiao et al. 2022) compared to $\mathcal{L}_{\textbf{InfoNCE}}$ and $\mathcal{L}_{\textbf{NC}}$. This suggests that downstream tasks can benefit from the learned representations obtained through our loss.
@@
To evaluate the performance of GCL-OT, the node classification tasks are conducted on 9 real-world datasets across different homophily degrees. Cora, PubMed (Sen et al. 2008), ArXiv, and Products (Hu et al. 2020) generally regarded as homophilic, whereas Amazon (Platonov et al. 2023), ArXiv23 (He et al. 2024), Wisconsin, Cornell, and Texas (Pei et al. 2020) fall into heterophilic category. Original textual information for heterophilic graphs was collected from publicly available sources (Leskovec et al. 2007; Craven et al. 1998; Wu et al. 2025).
Experiments are conducted on one NVIDIA 4090 GPU and three NVIDIA 3080 GPUs. Text augmentation uses GPT-3.5 following (He et al. 2024). The text encoder employs a partially frozen DistilBERT (Sanh et al. 2020) model with six layers. Graph encoder adopts GCN (Kipf and Welling 2017), GAT (Veličković et al. 2017), and GraphSAGE (Hamilton et al. 2017), each configured with two layers. GAT utilizes eight attention heads, and GraphSAGE samples all neighbors. The classifier adopts a three-layer MLP with ReLU. Dimension of Text and Graph encoder is set to 768, with learning rates of 0.01 and 2e-5, weight decay at 0.0005, dropout from 0 to 0.8, and a batch size of 512. ArXiv and Products follow the standard splits from (Hu et al. 2020), the other datasets use random splits (60/20/20) following (He et al. 2024). Baseline hyperparameters follow prior work. All experiments report the average accuracy and standard deviation over 10 random seeds.
@@
We first study the improvements over InfoNCE across various heterophily metrics by perturbing edges and texts. As shown in Figure 3, GCL-OT achieves substantial gains over the InfoNCE variant, obtained by replacing $\mathcal{L}_{\text{GCL-OT}}$ with $\mathcal{L}_{\text{InfoNCE}}$, with significant improvements under mixed-label neighborhoods and strong semantic heterophily. These trends indicate that GCL-OT can handle label mixing and semantic disagreement within neighborhoods, while the abundance of purely heterophilic neighborhoods and unlinked similar pairs may limit its benefits.
@@
(a) DistilBERT on text perturbation (%)
(b) GCL-OT on text perturbation (%)